# Chen and Peng

### v2

@@
Using the concept of Markov entanglement, we proved that a widely-used class of index policies is weakly entangled and enjoys a sublinear ${\mathcal{O}}(\sqrt{N})$ scale of decomposition error for $N$-agent systems. Finally, we show how Markov entanglement can be efficiently estimated in practice, providing practitioners with an empirical proxy for the quality of value decomposition.
@@ 1 Introduction
Learning the value function given certain policy, or policy evaluation, is one of the most fundamental tasks in RL. Significant attention has been paid to single-agent policy evaluation (Sutton and Barto 2018, Bertsekas and Tsitsiklis 1996, Tsitsiklis and Van Roy 1996). However, when it comes to multi-agent reinforcement learning (MARL), single-agent methodologies typically suffer from the curse of dimensionality: the state space of the system scales exponentially with the number of agents. To tackle this problem, one common technique is to decompose the global value function,
@@
In the past several decades, both positive and negative results have been reported. Back to the last century, Whittle (1988), Weber and Weiss (1990) apply Lagrange relaxations to decompose the global value and obtain the well-known Whittle index policy. The Lagrange decomposition idea has also been proved successful in many other important multi-agent tasks such as network revenue management (Adelman 2007, Zhang and Adelman 2009), resource allocation (Kadota et al. 2016, Balseiro et al. 2023), and online matching (Brown and Zhang 2022, 2023, Shar and Jiang 2023, Kanoria and Qian 2024). However, Lagrange decomposition relies on the knowledge of system dynamics and Adelman and Mersereau (2008) demonstrates its decomposition error can be arbitrarily bad for general weakly-coupled MDPs. In more recent days, practitioners apply online (deep) reinforcement learning to train a local value function for each individual agent. Albeit little theoretical understanding, this decomposition witnessed great empirical success in practice. For example, ride-hailing platforms conduct policy iteration with global value approximated by the summation of local value functions learned by individual drivers. This practice gives birth to state-of-the-art dispatching policies and have been well recognized by the operations research community, such as DiDi Chuxing (Qin et al. 2020, [Daniel H. Wagner Prize]) and Lyft (Azagirre et al. 2024, [Franz Edelman Laureates]). Intervention policies based on similar value decomposition idea also demonstrate substantial empirical advantage and have been deployed by a behavioral health platform in Kenya (Baek et al. 2023, [Pierskalla Award]). In broader MARL literature, value decomposition serves as one key component of centralized training and decentralized execution (CTDE) paradigm, achieving state-of-the-art performance (Sunehag et al. 2018, Rashid et al. 2020). In particular, agents optimize their local value functions and combining them to obtain the global optimal policy, adhering to the individual-global max (IGM) principle. However, recent research has started reflecting on invalidity and potential flaw of this principle in practice (Hong et al. 2022, Dou et al. 2022).
@@ 1.1 This Paper
where $\otimes$ is the tensor product or Kronecker product of matrices. The important question is whether we can extend beyond this trivial case of independent subsystems.
@@ A Sufficient and Necessary Condition
Compared with the preceding example of independent subsystems, Markov entanglement offers an intuitive interpretation: a two-agent MDP is separable if it can be expressed as a linear combination of independent subsystems. We then demonstrate,
separable $\textrm{separable }{\bm{P}}^{\pi}_{AB}\Longleftrightarrow\textrm{ decomposable}{\bm{V}}^{\pi}_{AB}\,,$
@@
where $\rho_{\rm{SEP}}$ is the set of all separable quantum states. In quantum physics, various distance measures have been designed for density matrices and capture different physical interpretations (Nielsen and Chuang 2010). In the Markov world, we analogously design distance measures for transition matrices and relate them to the value decomposition error,
@@
[…]
@@
We then leverage our Markov entanglement theory to analyze several structured multi-agent MDPs. We prove that a widely-used class of index policies is asymptotically separable, exhibiting a sublinear decomposition error that scales as ${\mathcal{O}}(\sqrt{N})$ with the number of agents $N$. This result theoretically justifies the practical effectiveness of value decomposition for index policies. Our proof builds on innovations that integrate Markov entanglement with mean-field analysis. Finally, we show that Markov entanglement admits an efficient empirical estimation, thus helping practitioners determine when value decomposition is feasible.
@@ 1.2 Other Related Work
Prior theoretical research has extensively investigated the decomposition of optimal value functions in multi-agent settings. A prominent area involves Lagrange relaxation, with the Restless Multi-Armed Bandit (RMAB, Whittle 1988) as a foundational model. The per-agent decomposition error is proven to decay asymptotically to zero (Weber and Weiss 1990, 1991, Verloop 2016), justifying the asymptotic optimality of the well-known Whittle Index policy (Whittle 1988). The decay rate is further refined to quadratic or exponential under various conditions (Gast et al. 2023, 2024, Brown and Zhang 2022, Zhang and Frazier 2021, 2022). Other work generalizes to Weakly-Coupled MDPs (WCMDPs), deriving guarantees based on system structure (Balseiro et al. 2021, Brown and Zhang 2025, Gast et al. 2022). However, Adelman and Mersereau (2008) showed Lagrange relaxation can have arbitrarily large errors and proposed an alternative decomposition called Approximate Linear Programs (ALP). ALP is proven to have tighter error, a finding further explored by Brown and Zhang (2023). Despite these advancements, characterizing decomposition error for general multi-agent MDPs remains challenging for both approaches. In contrast, our Markov entanglement theory analyzes value decomposition for general multi-agent MDPs under arbitrary policies, including optimal ones. Notably, we show sublinear decomposition error not only for the optimal Whittle Index policy but for any index policy as well.
Another line of theoretical work has concentrated on policy optimization via value decomposition. Despite reported empirical successes, rigorous theoretical analysis remains challenging. Baek et al. (2023) derived an approximation ratio for a specific index policy on a two-state RMAB. Wang et al. (2021), Dou et al. (2022) analyzed the convergence of the CTDE paradigm under strong exploration assumptions, while also highlighting scenarios of divergence. In contrast, our work instead focuses on policy evaluation rather than optimization. This enables us to derive clear and interpretable bounds on the decomposition error for general finite-state multi-agent MDPs that only require the existence of a stationary distribution.
Finally, we note that value decomposition can be viewed as single-agent policy evaluation with linear one-hot feature approximation. While extensive research has analyzed the convergence of single-agent policy evaluation with linear functions (Tsitsiklis and Van Roy 1996, Bhandari et al. 2021, Bertsekas and Tsitsiklis 1996, Srikant and Ying 2019, Liu and Olshevsky 2021), these results don’t explain how the limit point effectively approximates the global value—which is the central focus of our work. Thus, our paper addresses an aspect orthogonal to this literature.
@@ 1.3 Notations
We abbreviate subscripts $({\bm{s}})\coloneqq(s_{1:N})\coloneqq(s_{1},s_{2},\ldots,s_{N})$. Particularly, for two-agent case, when the context is clear, we abbreviate $({\bm{s}})\coloneqq(s_{AB})\coloneqq(s_{A},s_{B})$. Let $[N]=\{1,2,\ldots,N\}$ and $\mathbb{Z}^{+}$ be the set of positive integers. For a vector or matrix ${\bm{x}}$, ${\bm{x}}\geq 0$ is equivalent to every element of ${\bm{x}}$ to be non-negative and $\left\lvert{\bm{x}}\right\rvert$ denotes the element-wise absolute value. For (semi-)norm $\|\cdot\|_{\alpha}$ and norm $\|\cdot\|_{\beta}$, we define the $\alpha,\beta$-norm for matrix ${\bm{A}}$ as $\|{\bm{A}}\|_{\alpha,\beta}=\sup_{\|{\bm{x}}\|_{\beta}=1}\|{\bm{A}}{\bm{x}}\|_{\alpha}$. We further abbreviate $\|{\bm{A}}\|_{\alpha}\coloneqq\|{\bm{A}}\|_{\alpha,\alpha}$.
@@ 2 Model
where $\otimes$ is the tensor product and ${\bm{e}}=\mathbf{1}\in\mathbb{R}^{|S||A|}$ is the vector of all ones. This reward structure is broadly satisfied in various piratical scenarios. For example, a ride-hailing platform’s overall revenue is the summation of revenue of each driver. Many well-established models also include this reward structure, such as RMAB (Whittle 1988, Weber and Weiss 1990) and WCMDPs (Brown and Zhang 2022, Adelman and Mersereau 2008). In Appendix J.3, we extend our results to multi-agent MDP model where the global cannot be decomposed. We further assume the local rewards are bounded, i.e. for agent $i\in\{A,B\}$, $\left\lvert r_{i}(s_{i},a_{i})\right\rvert\leq r_{\max}^{i}$ for all $(s_{i},a_{i})$.
@@ 2.1 Local (Q-)value Functions
This meta-algorithm framework is simple and intuitive: each agent independently fits its local Q-values based on local observations. Notably, the framework requires no prior knowledge of the MDP, and learning can be performed in a fully decentralized manner. Furthermore, we use term meta in that we do not pose restrictions on how agents estimate their local Q-values. For tabular settings, one can plug in Temporal Difference (TD) learning (Sutton and Barto 2018) or its variants. For large-scale problems, one can apply linear function approximations (e.g. Baek et al. 2023, Han et al. 2022, Bertsekas and Tsitsiklis 1996) or more sophisticated neural networks (e.g. Qin et al. 2020, Sunehag et al. 2018, Mahajan et al. 2019).
@@ 3.1 An Illustrative Example of Coupling and Markov Entanglement
Consider a two-agent MDP ${\mathcal{M}}_{AB}$ with $\|{\mathcal{A}}_{A}\|=\|{\mathcal{A}}_{B}\|=2$ , where action $1$ means activate and 00 means idle. Each agent $i\in\{A,B\}$ has its own transition kernel ${\bm{P}}_{i}$. We examine the following policy:
@@ 3.2 Proof of Sufficiency
=$\overset{(i)}{=}\sum_{t=0}^{\infty}\gamma^{t}\left({\left({\sum_{j=1}^{K}x_{j}{\bm{P}}_{A}^{(j)}}\right)^{t}{\bm{r}}_{A}}\right)\otimes{\bm{e}}$
@@ 4 Necessary Condition for the Exact Value Decomposition
There exist other possible ways for value decomposition. For example, Sunehag et al. (2018), Dou et al. (2022) consider $Q^{\pi}_{AB}({\bm{s}},{\bm{a}})=L_{A}(s_{A},a_{A},{\bm{r}}_{AB})+L_{B}(s_{B},a_{B},{\bm{r}}_{AB})$ where $L_{A},L_{B}$ are learned jointly via minimizing the global Bellman error [^3] In Appendix E, we provide an example of entangled MDP that allows for such decomposition where $L_{A}$ depends on both ${\bm{r}}_{A}$ and ${\bm{r}}_{B}$. ; Rashid et al. (2020), Mahajan et al. (2019), Son et al. (2019), Wang et al. (2020) consider general monotonic operations beyond additive decompositions. These methods introduce possibly richer representations at the cost of more sophisticated implementations and less interpretability, which is beyond the scope of this paper.
@@ 4.1 Proof Sketch of Necessity
where $\varepsilon_{j}=(1,0,\ldots,0,-1,0,\ldots,0)^{\top}$ with the first element $1$ and $(j+1)$-th element $-1$.
@@
decomposable $\textrm{decomposable }Q^{\pi}_{AB}\Longrightarrow\textrm{ separable }(1-\gamma)({\bm{I}}-\gamma{\bm{P}}^{\pi}_{AB})^{-1}\,.$
@@
separable ($\textrm{separable }(1-\gamma)({\bm{I}}-\gamma{\bm{P}}^{\pi}_{AB})^{-1}\Longleftrightarrow\textrm{ separable }{\bm{P}}^{\pi}_{AB}\,.$
@@ 5 Value Decomposition Error in General Two-agent MDPs
where $\rho_{\rm{SEP}}$ is the set of all separable quantum states. In quantum physics, various distance measures have been developed for density matrices, including quantum relative entropy, Bures distance, and trace distance (Nielsen and Chuang 2010). These different distance measures embody distinct interpretations of the physical world.
@@ 5.2 Agent-wise Distance
The bound derived using ATV is more refined and explicitly takes into account the multi-agent structure. Each agent $i\in\{A,B\}$ contributes to the decomposition error by its entanglement with the system, i.e. $E_{i}({\bm{P}}^{\pi}_{AB})$, weighted by its maximum local reward $r^{i}_{\max}$. Furthermore, given that ATV is a tighter distance compared to TV, we have $E_{i}({\bm{P}}^{\pi}_{AB})\leq E({\bm{P}}^{\pi}_{AB})$ for $i\in\{A,B\}$, Theorem 5.5 provides a tighter bound for the entry-wise decomposition error compared to Theorem 5.3.
@@
In this section, we briefly work through the main proof framework and key techniques. The full proof is delayed to F. To begin, we can partition the decomposition error related to agent $A$ into two terms,
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}$ $\underbrace{({\bm{I}}-\gamma{\bm{P}}^{\pi}_{AB})^{-1}({\bm{r}}_{A}\otimes{\bm{e}})-\left({{\bm{I}}-\gamma{\bm{P}}_{A}\otimes{\bm{P}}_{B}}\right)^{-1}({\bm{r}}_{A}\otimes{\bm{e}})}_{(I)}+\underbrace{\left({\left({{\bm{I}}-\gamma{\bm{P}}_{A}}\right)^{-1}{\bm{r}}_{A}}\right)\otimes{\bm{e}}-\left({\left({{\bm{I}}-\gamma{\bm{P}}^{\pi}_{A}}\right)^{-1}{\bm{r}}_{A}}\right)\otimes{\bm{e}}}_{(II)}\,,$
@@
Lemma 5.6 (Lemma 1 in Farias et al. (2023))

### v3

@@
Using the concept of Markov entanglement, we prove that a widely-used class of index policies is weakly entangled and enjoys a sublinear ${\mathcal{O}}(\sqrt{N})$ scale of decomposition error for $N$-agent systems. Finally, we show how Markov entanglement can be efficiently estimated in practice, providing practitioners with an empirical proxy for the quality of value decomposition.
@@ 1 Introduction
Learning the value function given certain policy, or policy evaluation, is one of the most fundamental tasks in RL. Significant attention has been paid to single-agent policy evaluation (Sutton18RL, bertsekas1996neuro, john96analysis). However, when it comes to multi-agent reinforcement learning (MARL), single-agent methodologies typically suffer from the curse of dimensionality: the state space of the system scales exponentially with the number of agents. To tackle this problem, one common technique is to decompose the global value function,
@@
In the past several decades, both positive and negative results have been reported. Back to the last century, Whittle88restless, weber90on apply Lagrange relaxations to decompose the global value and obtain the well-known Whittle index policy. The Lagrange decomposition idea has also been proved successful in many other important multi-agent tasks such as network revenue management (adelman2007dynamic, zhang2009approximate), resource allocation (kadota2016minimizing, santiage23thebest), and online matching (brown2022dynamic, david23on, shar2023weakly, yash24blind). However, Lagrange decomposition relies on the knowledge of system dynamics and daniel08relaxation demonstrates its decomposition error can be arbitrarily bad for general weakly-coupled MDPs. In more recent days, practitioners apply online (deep) reinforcement learning to train a local value function for each individual agent. Albeit little theoretical understanding, this decomposition witnessed great empirical success in practice. For example, ride-hailing platforms conduct policy iteration with global value approximated by the summation of local value functions learned by individual drivers. This practice gives birth to state-of-the-art dispatching policies and have been well recognized by the operations research community, such as DiDi Chuxing (qin20ride, [Daniel H. Wagner Prize]) and Lyft (xabi24better, [Franz Edelman Laureates]). Intervention policies based on similar value decomposition idea also demonstrate substantial empirical advantage and have been deployed by a behavioral health platform in Kenya (baek2023policy, [Pierskalla Award]). In broader MARL literature, value decomposition serves as one key component of centralized training and decentralized execution (CTDE) paradigm, achieving state-of-the-art performance (peter18value, rashid2020monotonic). In particular, agents optimize their local value functions and combining them to obtain the global optimal policy, adhering to the individual-global max (IGM) principle. However, recent research has started reflecting on invalidity and potential flaw of this principle in practice (hong2022rethinking, dou2022understanding).
@@ 1.1 This Paper
where $\otimes$ is the tensor product or Kronecker product of matrices. The important question is whether we can extend beyond this trivial case of independent systems.
@@ A Sufficient and Necessary Condition
Compared with the preceding example of independent subsystems, Markov entanglement offers an intuitive interpretation: a two-agent MDP is separable if it can be expressed as a linear combination of independent systems. We then demonstrate,
$\textrm{separable }{\bm{P}}^{\pi}_{AB}\Longleftrightarrow\textrm{ decomposable }{\bm{V}}^{\pi}_{AB}\,$,
@@
where $\rho_{\rm{SEP}}$ is the set of all separable quantum states. In quantum physics, various distance measures have been designed for density matrices and capture different physical interpretations (nielsen2010quantum). In the Markov world, we analogously design distance measures for transition matrices and relate them to the value decomposition error,
@@
where $\left\lVert\cdot\right\rVert$ depends on the distance we use to measure Markov entanglement. We explore diverse distance measures including the well-known total variation distance and its stationary distribution weighted variant. We also design a novel agent-wise distance incorporating the multi-agent structure, which may be of independent interest to the MARL community. We further demonstrate how different distance measures capture the entanglement from different perspectives and give birth to the decomposition error in different norms.
@@
We then apply Markov entanglement theory to several structured multi-agent MDPs. We prove a widely-used class of index policies is asymptotically separable, exhibiting a sublinear decomposition error scaling as ${\mathcal{O}}(\sqrt{N})$ with the number of agents $N$. This result theoretically justifies the practical effectiveness of value decomposition for index policies. Our proof builds on innovations that integrate Markov entanglement with mean-field analysis.
For practitioners, we show that Markov entanglement can be efficiently estimated and serves as a surrogate to test whether value decomposition is feasible. Finally, we empirically demonstrate the low-entangled structure in several practical scenarios including a ride-hailing simulator.
@@ 1.2 Other Related Work
Prior theoretical research has extensively investigated the decomposition of optimal value functions in multi-agent settings. A prominent area involves Lagrange relaxation, with the Restless Multi-Armed Bandit (RMAB, Whittle88restless) as a foundational model. The per-agent decomposition error is proven to decay asymptotically to zero (weber90on, Weber_Weiss_1991, Verloop2016AsymptoticallyOP), justifying the asymptotic optimality of the well-known Whittle Index policy (Whittle88restless). The decay rate is further refined to quadratic or exponential under various conditions (gast23exponential, gast24linear, brown2022dynamic, zhang2021restless, zhang2022near). Other work generalizes to Weakly-Coupled MDPs (WCMDPs), deriving guarantees based on system structure (santiago21, david25fluid, gast2022reoptimization). However, daniel08relaxation showed Lagrange relaxation can have arbitrarily large errors and proposed an alternative decomposition called Approximate Linear Programs (ALP). ALP is proven to have tighter error, a finding further explored by david23on. Despite these advancements, characterizing decomposition error for general multi-agent MDPs remains challenging for both approaches. In contrast, our Markov entanglement theory analyzes value decomposition for general multi-agent MDPs under arbitrary policies, including optimal ones. Notably, we show sublinear decomposition error not only for the optimal Whittle Index policy but for any index policy as well.
Another line of theoretical work has concentrated on policy optimization via value decomposition. Despite the reported empirical successes, rigorous theoretical analysis remains challenging. baek2023policy derived an approximation ratio for a specific index policy on a two-state RMAB. wang2021towards, dou2022understanding analyzed the convergence of the CTDE paradigm under strong exploration assumptions, while also highlighting scenarios of divergence. In contrast, our work instead focuses on policy evaluation rather than optimization. This enables us to derive clear and interpretable bounds on the decomposition error for general finite-state multi-agent MDPs that only require the existence of a stationary distribution.
Finally, we note that value decomposition can be viewed as single-agent policy evaluation with linear one-hot feature approximation. While extensive research has analyzed the convergence of single-agent policy evaluation with linear functions (john96analysis, jalaj21afinite, bertsekas1996neuro, srikant2019finite, liu2021temporal), these results don’t explain how the limit point effectively approximates the global value—which is the central focus of our work. Thus, our paper addresses an aspect orthogonal to this literature.
@@ 1.3 Notations
We abbreviate subscripts $({\bm{s}})\coloneqq(s_{1:N})\coloneqq(s_{1},s_{2},\ldots,s_{N})$. Particularly, for two-agent case, when the context is clear, we abbreviate $({\bm{s}})\coloneqq(s_{AB})\coloneqq(s_{A},s_{B})$. Let $N]=\{1,2,\ldots,N\}$ and $\mathbb{Z}^{+}$ be the set of positive integers. For a vector or matrix $\bm{x}$}, ${\bm{x}}\geq 0$ is equivalent to every element of $\bm{x}$} to be non-negative and $\left\lvert{\bm{x}}\right\rvert$ denotes the element-wise absolute value. For (semi-)norm $\|\cdot\|_{\alpha}$ and norm $\|\cdot\|_{\beta}$, we define the $\alpha,\beta$-norm for matrix $\bm{A}$} as $\|{\bm{A}}\|_{\alpha,\beta}=\sup_{\|{\bm{x}}\|_{\beta}=1}\|{\bm{A}}{\bm{x}}\|_{\alpha}$. We further abbreviate $\|{\bm{A}}\|_{\alpha}\coloneqq\|{\bm{A}}\|_{\alpha,\alpha}$.
@@ 2 Model
where $\otimes$ is the tensor product and ${\bm{e}}=\mathbf{1}\in\mathbb{R}^{|S||A|}$ is the vector of all ones. This reward structure is broadly satisfied in various piratical scenarios. For example, a ride-hailing platform’s overall revenue is the summation of revenue of each driver. Many well-established models also include this reward structure, such as RMAB (Whittle88restless, weber90on) and WCMDPs (brown2022dynamic, daniel08relaxation). In Appendix J.3, we extend our results to multi-agent MDP model where the global cannot be decomposed. We further assume the local rewards are bounded, i.e. for agent $i\in\{A,B\}$, $\left\lvert r_{i}(s_{i},a_{i})\right\rvert\leq r_{\max}^{i}$ for all $(s_{i},a_{i})$.
@@ 2.1 Local (Q-)value Functions
This meta-algorithm framework is simple and intuitive: each agent independently fits its local Q-values based on local observations. Notably, the framework requires no prior knowledge of the MDP, and learning can be performed in a fully decentralized manner. Furthermore, we use term meta in that we do not pose restrictions on how agents estimate their local Q-values. For tabular settings, one can plug in Temporal Difference (TD) learning (Sutton18RL) or its variants. For large-scale problems, one can apply linear function approximations (e.g. baek2023policy, han22real, bertsekas1996neuro) or more sophisticated neural networks (e.g. qin20ride, peter18value, mahajan2019maven).
@@ 3.1 An Illustrative Example of Coupling and Markov Entanglement
Consider a two-agent MDP ${\mathcal{M}}_{AB}$ with $\|{\mathcal{A}}_{A}\|=\|{\mathcal{A}}_{B}\|=2$ , where action $1$ means activate and 0 means idle. Each agent $i\in\{A,B\}$ has its own transition kernel ${\bm{P}}_{i}$. We examine the following policy:
@@ 3.2 Proof of Sufficiency
$\overset{(i)}{=}\sum_{t=0}^{\infty}\gamma^{t}\left({\left({\sum_{j=1}^{K}x_{j}{\bm{P}}_{A}^{(j)}}\right)^{t}{\bm{r}}_{A}}\right)\otimes{\bm{e}}$
@@ 4 Necessary Condition for the Exact Value Decomposition
There exist other possible ways for value decomposition. For example, peter18value, dou2022understanding consider $Q^{\pi}_{AB}({\bm{s}},{\bm{a}})=L_{A}(s_{A},a_{A},{\bm{r}}_{AB})+L_{B}(s_{B},a_{B},{\bm{r}}_{AB})$ where $L_{A},L_{B}$ are learned jointly via minimizing the global Bellman error [^3] In Appendix E, we provide an example of entangled MDP that allows for such decomposition where $L_{A}$ depends on both ${\bm{r}}_{A}$ and ${\bm{r}}_{B}$. ; rashid2020monotonic, mahajan2019maven, son19qtran, wang2020qplex consider general monotonic operations beyond additive decompositions. These methods introduce possibly richer representations at the cost of more sophisticated implementations and less interpretability, which is beyond the scope of this paper.
@@ 4.1 Proof Sketch of Necessity
where $\varepsilon_{j}=(1,0,\ldots,0,-1,0,\ldots,0)^{\top}$ with the first element $1$ and $(j+1)$-th element 1-1.
@@
$\textrm{decomposable }Q^{\pi}_{AB}\Longrightarrow\textrm{ separable }(1-\gamma)({\bm{I}}-\gamma{\bm{P}}^{\pi}_{AB})^{-1}\,.$
@@
$\textrm{separable }(1-\gamma)({\bm{I}}-\gamma{\bm{P}}^{\pi}_{AB})^{-1}\Longleftrightarrow\textrm{ separable }{\bm{P}}^{\pi}_{AB}\,.$
@@ 5 Value Decomposition Error in General Two-agent MDPs
where $\rho_{\rm{SEP}}$ is the set of all separable quantum states. In quantum physics, various distance measures have been developed for density matrices, including quantum relative entropy, Bures distance, and trace distance (nielsen2010quantum). These different distance measures embody distinct interpretations of the physical world.
@@ 5.2 Agent-wise Distance
The bound derived using ATV is more refined and explicitly takes into account the multi-agent structure. Each agent $i\in\{A,B\}$ contributes to the decomposition error by its entanglement with the system, i.e. $E_{i}({\bm{P}}^{\pi}_{AB})$, weighted by its maximum local reward $r^{i}_{\max}$. Furthermore, given that ATV is a tighter distance compared to TV, i.e. $E_{i}({\bm{P}}^{\pi}_{AB})\leq E({\bm{P}}^{\pi}_{AB})$ for $i\in\{A,B\}$, Theorem 5.5 provides a tighter bound for the entry-wise decomposition error compared to Theorem 5.3.
@@
In this section, we briefly work through the main proof framework and key techniques. The full proof is delayed to Appendix F. To begin, we can partition the decomposition error related to agent $A$ into two terms,
@@
$\stackrel{{\scriptstyle(i)}}{{=}}$ $\underbrace{({\bm{I}}-\gamma{\bm{P}}^{\pi}_{AB})^{-1}({\bm{r}}_{A}\otimes{\bm{e}})-\left({{\bm{I}}-\gamma{\bm{P}}_{A}\otimes{\bm{P}}_{B}}\right)^{-1}({\bm{r}}_{A}\otimes{\bm{e}})}_{(I)}+\underbrace{\left({\left({{\bm{I}}-\gamma{\bm{P}}_{A}}\right)^{-1}{\bm{r}}_{A}}\right)\otimes{\bm{e}}-\left({\left({{\bm{I}}-\gamma{\bm{P}}^{\pi}_{A}}\right)^{-1}{\bm{r}}_{A}}\right)\otimes{\bm{e}}}_{(II)}\,$,
@@
Lemma 5.6 (Lemma 1 in farias2023correcting)